# ClustRecNet: A Novel End-to-End Deep Learning Framework for Clustering Algorithm Recommendation

### v2

@@ Abstract
We introduce ClustRecNet – a novel deep learning (DL)-based recommendation framework for determining the most suitable clustering algorithms for a given dataset, addressing the long-standing challenge of clustering algorithm selection in unsupervised learning. To enable supervised learning in this context, we construct a comprehensive data repository comprising 34,000 synthetic datasets with diverse structural properties. Each of them was processed using 10 popular clustering algorithms. The resulting clusterings were assessed via the Adjusted Rand Index (ARI) to establish ground truth labels, used for training and evaluation of our DL model.
The proposed network architecture integrates convolutional, residual, and attention mechanisms to capture both local and global structural patterns from the input data. This design supports end-to-end training to learn compact representations of datasets and enables direct recommendation of the most suitable clustering algorithm, reducing reliance on handcrafted meta-features and traditional Cluster Validity Indices (CVIs). Comprehensive experiments across synthetic and real-world benchmarks demonstrate that our DL model consistently outperforms conventional CVIs (e.g. Silhouette, Calinski-Harabasz, Davies-Bouldin, and Dunn) as well as state-of-the-art AutoML clustering recommendation approaches (e.g. ML2DAC, AutoCluster, and AutoML4Clust). Notably, the proposed model achieves a 0.497 ARI improvement over the Calinski-Harabasz index on synthetic data and a 15.3% ARI gain over the best-performing AutoML approach on real-world data.
Codehttps://github.com/confanonymgit/clustrecnet
@@ 1 Introduction
Selecting an appropriate clustering algorithm for an unlabeled dataset remains one of the most elusive tasks in unsupervised learning. Due to the diversity of input data, numerous clustering algorithms have been proposed in the literature, each possessing its strengths and limitations. Several studies have been dedicated to the comparison and categorization of clustering algorithms (Brock et al. 2008; Arbelaitz et al. 2013; Rodriguez et al. 2019; Wani 2024), providing valuable insights into their advantages and limitations.
@@
Despite these significant contributions, selecting the most suitable clustering algorithm for a given dataset remains a complex task, necessitating considerable domain knowledge and extensive trial-and-error experimentation. A crucial step in this process involves evaluating the quality of clustering solutions, which is typically carried out using established cluster validity indices (CVIs) (Hassan et al. 2024). Indices such as Silhouette (Sil) (Rousseeuw 1987), Calinski-Harabasz (CH) (Caliński and Harabasz 1974), Davies–Bouldin (DB) (Davies and Bouldin 1979), and Dunn (Dunn 1973) reflect several criteria of cluster compactness and separation. However, these CVIs often fall short when dealing with datasets characterized by complex high-dimensional structures, outliers, or overlapping clusters, thereby revealing some important limitations in existing cluster validation methodologies.
@@
Recently, several AutoML-based approaches have been introduced to address the challenge of selecting the most suitable clustering algorithm for a given dataset in unsupervised settings (Francia, Giovanelli, and Golfarelli 2024). Many of them are inspired by established AutoML techniques in supervised learning, particularly those tackling the Combined Algorithm Selection and Hyperparameter optimization (CASH) problem (Thornton et al. 2013). The literature predominantly outlines two main strategies. The first of them jointly optimizes the algorithm choice and its hyperparameter space, as exemplified by AutoML4Clust (Tschechlov et al. 2021). However, such an approach is often computationally expensive, particularly on large and complex search spaces. The second, more prevalent, strategy treats the CASH problem sequentially — first selecting a suitable algorithm, followed by hyperparameter tuning. The approaches supporting this strategy leverage meta-learning techniques that transfer knowledge from existing datasets.
@@
Early work by De Souto et al. (2008) pioneered the use of meta-learning for selecting the most appropriate clustering algorithm(s), thus laying the groundwork for subsequent research in this domain. Most AutoML systems for clustering follow a two-stage architecture: (1) Offline phase that extracts fixed-length meta-features and associates them with algorithm performance; (2) Online phase that employs shallow learners, such as $k$-nearest neighbors or decision trees, to recommend the most convenient clustering algorithms for new datasets. For example, AutoCluster (Liu, Li, and Tian 2021) leverages landmark-based and statistical meta-features, applying a grid search strategy for configuration optimization. In contrast, ML2DAC (Treder-Tschechlov et al. 2023) relies on a diversified set of data descriptors, including general properties (e.g. instance counts), statistical moments, and information-theoretic measures, to recommend both clustering algorithms and their optimal hyperparameters.
@@
Despite these contributions, existing meta-learning-based methods face several inherent limitations. The reliance on fixed-length meta-feature vectors introduces an abstraction layer that may obscure essential structural and distributional characteristics of the data, which critically influence clustering behavior and algorithm performance. Moreover, the selection of informative and representative meta-features remains an open problem, with no established consensus on optimal descriptors (Poulakis, Doulkeridis, and Kyriazis 2024). Finally, the prevailing modular architecture – separating feature extraction from model training – limits the expressiveness and adaptability of learned representations, potentially constraining generalization to complex datasets.
In this paper, we propose a novel deep learning (DL) framework that eliminates the need for handcrafted meta-feature engineering or proxy representations. By treating each dataset holistically as a learning instance, we train a deep neural network in an end-to-end fashion to learn a direct mapping from raw data to algorithm recommendation. This design enables the model to capture high-level structural patterns directly from the underlying data distribution. As a result, it delivers more accurate and generalizable clustering algorithm recommendations across a diverse collection of synthetic and real-world datasets.
@@
The paper is organized as follows: Section 2 briefly presents the data generation process, clustering algorithms being used, the proposed ClustRecNet model architecture, and the training process. Section 3 describes the experimental setup and discusses the results obtained from experiments with synthetic and real data. Finally, Section 4 concludes the study, summarizing its key findings and implications.
@@ 2 Methodology
The proposed recommendation system introduces a novel framework for selecting clustering algorithms, which overcomes key limitations of traditional validation indices and recent AutoML approaches. Our method employs a DL model trained on a diverse repository of 34,000 datasets, enabling it to learn structural patterns across varying clustering scenarios. To generate training labels, we evaluate the performance of 10 widely-used clustering algorithms on each dataset considered, using the Adjusted Rand Index (ARI) (Hubert and Arabie 1985) (i.e. the best-performing algorithm(s) correspond to the highest ARI value(s)). Once trained on these ARI-labeled datasets, our model can effectively recommend the most suitable clustering algorithms for previously unseen data. This design not only automates the selection process but also improves clustering accuracy and generalization, eliminating the need for handcrafted meta-features and CVIs.
@@ 2.1 Data Generation
To train a robust DL-based recommendation framework, we generate synthetic datasets with controlled variation across key parameters, including the number and shape of clusters, the number of objects and features, intra-cluster density, degree of cluster overlap, and level of noise added to feature values. Accounting for these factors ensures that the model can generalize across a wide range of clustering scenarios and provide accurate algorithm recommendations.
@@
1) Scenario 1 follows the data generation strategy proposed by Rodriguez et al. (2019). It relies on five key parameters: $K$ - the number of clusters in the dataset, $N$ - the number of objects, $D$ - the number of features (i.e. number of dimensions) for each object, $N_{e}$ - the number of objects per cluster, and $\alpha$ - the cluster separation parameter (i.e. degree of overlap between clusters). To cover a broad range of clustering complexities in our simulations, we varied the number of clusters, features, and objects per cluster. Specifically, we considered the values of 2, 5, and 10 for the number of clusters; 2, 5, 10, and 50 for the number of features (to simulate both low- and high-dimensional data); and 50 and 200 for the number of objects per cluster. This corresponds to 24 different parameter configurations. To select appropriate values of the $\alpha$ parameter, which controls inter-cluster separation and depends on both cluster count and dimensionality, we performed a grid search over the range [0.1, 10] with a step size of 0.1. The selected $\alpha$ values satisfied two criteria: No single clustering algorithm consistently achieved an ARI above 0.8 across all datasets of any given parameter configuration, and at least one algorithm exceeded this threshold in more than 20% of the realizations. These constraints prevented any single algorithm from clearly dominating the others. In total, we identified 11 parameter configurations meeting these criteria.
@@
2) Scenario 2 corresponds to an alternative data generation strategy proposed by Zellinger and Bühlmann (2023). We used this strategy when no values of $\alpha$ were satisfying the desired ARI conditions for a given parameter configuration. Instead of considering the cluster separation parameter $\alpha$, the procedure of Zellinger and Buhlmann relies on a richer set of parameters, including cluster overlap, aspect ratio, radius ratio, cluster distribution, and imbalance ratio. In these settings, we explored three different minimum-maximum overlap intervals: $(0.001,0.002)$, $(0.001,0.25)$, and $(0.25,0.30)$. As the aspect ratio (i.e. the ratio between the longest and the shortest axis in a cluster), we used the values of 1 and 3. As the radius ratio (i.e. the ratio between the radii of the largest and the smallest clusters), we used the values of 1, 3, and 10. Cluster distributions considered included normal, exponential, and $t$ distributions. Finally, the cluster imbalance ratios of 1, 3, and 5 were explored. To mitigate bias, we calibrated parameter combinations under the same conditions as in Scenario 1. As a result, we identified 6 different valid parameter configurations.
Figure 1: Scatter plots illustrating the distribution of data points in two sample datasets: (a, Scenario 1) Clusters generated according to the strategy of Rodriguez et al. (2019), using the following parameters: $K=5$, $N=1000$, $D=2$, $N_{e}=200$, and $\alpha=5$; (b, Scenario 2) Clusters generated according to the strategy of Zellinger and Bühlmann (2023), using the following parameters: $K=10$, $N=500$, $D=2$, cluster overlap $(0.001,0.002)$, aspect ratio 33, radius ratio 33, normal distribution, and imbalance ratio 55.
@@
Overall, we generated 2,000 different datasets for each of the 17 retained configurations (11 from Scenario 1 and 6 from Scenario 2) to ensure an empirical trade-off between dataset diversity and training stability, resulting in a comprehensive data repository of 34,000 synthetic datasets (available at our GitHub). Figure 1 illustrates examples of datasets considered in Scenarios 1 and 2. To ensure a consistent input scale for DL training, we normalized all datasets using $z$-scores. Additionally, we applied symmetric padding to align input shapes across varying dataset sizes.
@@ 2.2 Clustering Algorithms
The set of clustering algorithms considered in our study include partitioning methods such as $k$-means (McQueen 1967) and $k$-medians (Jain and Dubes 1988), hierarchical approaches such as Agglomerative Clustering (AC) (Defays 1977), BIRCH (Zhang, Ramakrishnan, and Livny 1996), and Ward’s method (Ward Jr 1963), and density-based techniques such as DBSCAN (Ester et al. 1996), HDBSCAN (Campello, Moulavi, and Sander 2013), and OPTICS (Ankerst et al. 1999). Moreover, we used the probabilistic Gaussian Mixture Models (GMM) (Reynolds et al. 2009) and the graph-based Spectral Clustering (SC) (Ng, Jordan, and Weiss 2001) clustering techniques for their unique properties and advantages. All implementations were carried out using the scikit-learn library, and where required (e.g. in $k$-means), the number of clusters was assumed to be known. Given the extensive literature on parameter selection for clustering algorithms, we did not analyze parameter effects comprehensively. Therefore, instead of optimizing hyperparameters for each realization, we performed grid search once per configuration to maximize the mean of ARI across all 2,000 datasets generated under each of our 17 parameter configurations, thus ensuring consistent and reproducible tuning, while significantly reducing computational overhead. The hyperparameter selection details for the 10 clustering algorithms considered in our study are available in Supplementary Material.
@@ 2.3 Deep Model Architecture
It starts with a convolutional layer, followed by batch normalization and max pooling. Integrating the 2D convolutional layer, batch normalization, and max pooling is a standard and effective way of organizing an initial processing block in contemporary CNN architectures. After the convolutional layer, the architecture incorporates two consecutive residual blocks. Each of them can be characterized by the following equation:
@@
After the residual blocks, we implement a self-attention mechanism inspired by the transformer architecture from Vaswani et al. (2017). The attention block can be described by the following equation:
@@
where $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ are learned linear projections of the input features. Specifically $\mathbf{Q}=\text{Conv2D}(\mathbf{W}_{q},\mathbf{X}),\mathbf{K}=\text{Conv2D}(\mathbf{W}_{k}\mathbf{X})$, and $\mathbf{V}=\text{Conv2D}(\mathbf{W}_{v},\mathbf{X})$, where $\mathbf{W}_{q}$, $\mathbf{W}_{k}$, and $\mathbf{W}_{v}$ are the respective learned weight matrices. The attention block uses the learned query, key, and value projections to compute a weighted sum of the input features, enabling the model to capture long-range dependencies within the feature maps and complement the local processing capabilities of the convolutional layers. The output from the attention block is subsequently flattened and processed through two fully connected layers, reaching the final layer that produces logits corresponding to the number of classes in the classification task.
Therefore, the proposed network architecture is designed to integrate local spatial pattern recognition (via CNNs), stable and hierarchical feature propagation (via residual connections), and global context modeling (via attention), offering a unified solution for learning discriminative representations from raw tabular data. As we will see in the next sections, an effective combination of these components allows our ClustRecNet model to deliver quality performance on complex clustering classification tasks.

### v3

@@ Abstract
In unsupervised learning, identifying an effective clustering algorithm for a given tabular dataset remains a fundamental challenge. We introduce ClustRecNet, a novel end-to-end deep learning framework that recommends a suitable clustering algorithm by directly learning high-order representations of raw tabular data. To facilitate robust meta-learning, we construct a comprehensive repository of $34,000$ synthetic datasets with diverse structures, run 10 prominent clustering algorithms, and use Adjusted Rand Index (ARI) to establish ground-truth labels. ClustRecNet integrates convolutional, residual, and attention mechanisms to capture both local/global structural patterns, effectively bypassing the knowledge bottleneck associated with manual feature engineering. Extensive evaluations on both synthetic and real-world benchmarks demonstrate that ClustRecNet consistently outperforms state-of-the-art Automated Machine Learning (AutoML) approaches, including ML2DAC and AutoML4Clust. Our framework achieves an average $0.497$ ARI gain over the well-known Calinski–Harabasz cluster validity index on synthetic data and an average $15.3\%$ ARI improvement over the leading AutoML approach (ML2DAC) on real-world benchmarks. To the best of our knowledge, we are the first to successively apply deep learning to automatically recommend suitable clustering algorithms for tabular data at hand.
keywords:
Clustering , Unsupervised Learning , Deep Learning , Automated Machine Learning (AutoML) , Adjusted Rand Index (ARI) , Cluster Validity Indices (CVIs)
[…]
[4] organization=Department of Computer Science and Operations Research, Université de Montréal, city=Montreal, state=QC, country=Canada
@@ 1 Introduction
Clustering is a cornerstone of knowledge discovery, enabling the extraction of patterns from unlabeled data in complex domains such as medical informatics, social network analysis, and anomaly detection [20, 13, 22, 36]. However, the utility of the extracted knowledge is strictly influenced by the selection of an appropriate clustering algorithm. Despite the proliferation of diverse clustering methodologies, each characterized by specific inductive biases and operational constraints, identifying a suitable clustering algorithm for a given dataset remains a non-trivial challenge. While existing comparative frameworks [3, 2, 25, 32] offer valuable insights, the lack of an automated, data-driven mechanism for clustering algorithm selection remains a significant bottleneck in the development of fully autonomous intelligent systems.
@@
This challenge is further compounded by the difficulty of objectively validating clustering outcomes in the absence of prior labels, necessitating considerable domain knowledge and extensive trial-and-error experimentation. A crucial step in this process involves evaluating the quality of clustering solutions, which is typically carried out using established cluster validity indices (CVIs) [15, 7, 27]. The indices such as Silhouette (Sil) [26], Calinski-Harabasz (CH) [4], Davies–Bouldin (DB) [6], and Dunn [11] reflect different criteria of cluster compactness and separation. However, these CVIs often fall short when dealing with datasets characterized by complex structures, outliers, or overlapping clusters, thereby revealing critical limitations in existing cluster validation methodologies.
@@
Recently, several Automated Machine Learning (AutoML) approaches have been introduced to address the issue of selecting the most suitable clustering algorithm for a given dataset in unsupervised settings [14, 10]. Many of them are inspired by established AutoML techniques in supervised learning, particularly those tackling the Combined Algorithm Selection and Hyperparameter optimization (CASH) problem [28]. The literature predominantly outlines two main strategies. The first of them jointly optimizes the choice of clustering algorithm and its hyperparameter space, as exemplified by AutoML4Clust [30]. However, such an approach is often computationally expensive, particularly on large and complex search spaces. The second, more prevalent, strategy treats the CASH problem sequentially — first selecting a suitable clustering algorithm, followed by hyperparameter tuning. The approaches supporting this strategy leverage meta-learning techniques that transfer knowledge from existing datasets.
@@
Early work by De Souto et al. [8] pioneered the use of meta-learning for selecting the most appropriate clustering algorithm(s) for a given dataset, thus laying the groundwork for subsequent research in this domain. Most AutoML systems for clustering follow a two-stage architecture: (1) an offline phase that extracts fixed-length meta-features and associates them with algorithm performance; (2) an online phase that employs shallow learners, such as $k$-nearest neighbors or decision trees, to recommend the most convenient clustering algorithms for new datasets. For example, AutoCluster [18] leverages landmark-based and statistical meta-features, applying a grid search strategy for configuration optimization. In contrast, ML2DAC [29] relies on a diversified set of data descriptors, including general properties (e.g. instance counts), statistical moments, and information-theoretic measures, to recommend both clustering algorithms and their optimal hyperparameters.
@@
Despite their key strengths, existing meta-learning-based methods face several inherent limitations. The reliance on manually engineered meta-feature vectors introduces a representation gap that may obscure essential structural knowledge and distributional characteristics of the data, which critically influence clustering behavior and algorithm performance. Moreover, the selection of informative and representative meta-features remains an open problem, with no established consensus on optimal descriptors [23]. Finally, the prevailing modular architecture – separating feature extraction from model training – limits the expressiveness and adaptability of learned representations, potentially constraining generalization to complex datasets.
In this paper, we propose ClustRecNet, a novel deep learning framework that eliminates the need for handcrafted meta-feature engineering and facilitates the automated discovery of data geometry to recommend suitable clustering algorithms for tabular data at hand. The main contributions of our work are as follows:
  • 1.
    We introduce an end-to-end deep learning architecture that learns latent structural representations directly from input data distributions to recommend suitable clustering algorithms.
  • 2.
    We eliminate the information bottleneck caused by manual meta-feature engineering, allowing the model to capture high-order structural patterns.
  • 3.
    We carry out an extensive empirical evaluation of the new model on both synthetic and real-world benchmarks, demonstrating its superior generalization capability compared to three state-of-the-art AutoML methods (AutoCluster, AML4C, and ML2DAC) and four traditional cluster validity indices (Silhouette, Calinski-Harabasz, Davies–Bouldin, and Dunn).
@@
The paper is organized as follows: Section 2 presents the data generation protocol, clustering algorithms being used, the proposed ClustRecNet architecture, and the training process. Section 3 describes the experimental setup and reports results from experiments with synthetic and real data. Finally, Section 4 concludes the study, summarizing its key findings and implications.
@@ 2 Methodology
The proposed recommendation system introduces a novel deep learning framework for selecting clustering algorithms, which overcomes key limitations of traditional validation indices and recent AutoML approaches. Our method employs a network architecture trained on a diverse repository of 34,000 datasets, enabling it to learn structural patterns across varying clustering scenarios. To generate training labels, we evaluated the performance of 10 widely-used clustering algorithms on each dataset considered, using the Adjusted Rand Index (ARI) [16] (i.e. the best-performing algorithm(s) correspond to the highest ARI value(s)). Once trained on these ARI-labeled datasets, our model can effectively recommend the most suitable clustering algorithms for previously unseen data. As we will see, such an ARI-based training not only automates the clustering algorithm selection process but also improves clustering accuracy and generalization, eliminating the need for handcrafted meta-features and CVIs.
@@ 2.1 Data Generation
To train a robust DL (deep learning)-based recommendation framework, we generated synthetic datasets with controlled variation across key parameters, including the number and shape of clusters, the number of objects and features, intra-cluster density, degree of cluster overlap, and level of noise added to feature values. Accounting for these factors ensures that the model can generalize across a wide range of clustering scenarios and provide accurate clustering algorithm recommendations.
@@
1) Scenario 1 follows the data generation strategy proposed by Rodriguez et al. [25]. It relies on five key parameters: $K$ - the number of clusters in the dataset, $N$ - the number of objects in the dataset, $D$ - the number of features (i.e. number of available dimensions) for each object, $N_{e}$ - the number of objects per cluster, and $\alpha$ - the cluster separation parameter (i.e. degree of overlap between clusters). To cover a broad range of clustering complexities in our simulations, we varied the number of clusters, features, and objects per cluster. Specifically, we considered the values of 2, 5, and 10 for the number of clusters; 2, 5, 10, and 50 for the number of features; and 50 and 200 for the number of objects per cluster. This corresponds to 24 different parameter configurations. To select appropriate values of the $\alpha$ parameter, which controls inter-cluster separation and depends on both the number of clusters and dimensionality, we performed a grid search over the range [0.1, 10] with a step size of 0.1. The selected $\alpha$ values satisfied two criteria: (1) no single clustering algorithm consistently achieved an ARI above 0.8 across all datasets of any given parameter configuration, and (2) at least one algorithm exceeded this threshold in more than 20% of the realizations. These constraints prevented any single algorithm from clearly dominating the others. In total, we identified 11 parameter configurations meeting these criteria.
@@
2) Scenario 2 corresponds to an alternative data generation strategy proposed by Zellinger and Buhlmann [34]. We used this strategy when no values of $\alpha$ were satisfying the desired ARI conditions for a given parameter configuration. Instead of considering the cluster separation parameter $\alpha$, the procedure of Zellinger and Buhlmann relies on a richer set of parameters, including cluster overlap, aspect ratio, radius ratio, cluster distribution, and imbalance ratio. In these settings, we explored three different minimum-maximum overlap intervals: $(0.001,0.002)$, $(0.001,0.25)$, and $(0.25,0.30)$. As the aspect ratio (i.e. the ratio between the longest and the shortest axes in a cluster), we used the values of 1 and 3. As the radius ratio (i.e. the ratio between the radii of the largest and the smallest clusters), we used the values of 1, 3, and 10. Cluster distributions considered included normal, exponential, and $t$ distributions. Finally, the cluster imbalance ratios of 1, 3, and 5 were explored. To mitigate bias, we calibrated parameter combinations under the same conditions as in Scenario 1. As a result, we identified 6 different valid parameter configurations for this scenario.
@@
Overall, we generated 2,000 datasets for each of the 17 retained configurations (11 from Scenario 1 and 6 from Scenario 2) to ensure an empirical trade-off between dataset diversity and training stability, resulting in a comprehensive data repository of 34,000 synthetic datasets (available at our GitHub repository [^1] https://github.com/mrbakhtyari/ClustRecNet). Figure 1 presents examples of datasets from Scenarios 1 and 2. To ensure a consistent input scale for training, we normalized the features in all datasets using $z$-scores. Additionally, we applied symmetric padding to align the input shapes across datasets of varying sizes.
Figure 1: Scatter plots illustrating the distribution of objects in two sample datasets: (a, Scenario 1) Clusters generated according to the strategy of Rodriguez et al. [25], using the following parameters: $K=5$, $N=1000$, $D=2$, $N_{e}=200$, and $\alpha=5$; (b, Scenario 2) Clusters generated according to the strategy of Zellinger and Buhlmann [34], using the following parameters: $K=10$, $N=500$, $D=2$, cluster overlap $(0.001,0.002)$, aspect ratio 33, radius ratio 33, normal distribution, and imbalance ratio 55.
@@ 2.2 Clustering Algorithms
The set of clustering algorithms considered in our study include partitioning methods such as $k$-means [19] and $k$-medians [17], hierarchical approaches such as Agglomerative Clustering (AC) [9], BIRCH [35], and Ward’s method [33], and density-based techniques such as DBSCAN [12], HDBSCAN [5], and OPTICS [1]. Moreover, we used the probabilistic Gaussian Mixture Models (GMM) [24] and the graph-based Spectral Clustering (SC) [21] techniques for their unique properties and advantages. All implementations were carried out using the scikit-learn library, and where required (e.g. in $k$-means), the number of clusters was assumed to be known. Given the extensive literature on parameter selection for clustering algorithms, we did not analyze parameter effects comprehensively. Therefore, instead of optimizing hyperparameters for each realization, we performed grid search once per configuration to maximize the mean of ARI across all 2,000 datasets generated under each of our 17 parameter configurations, thus ensuring consistent and reproducible tuning while significantly reducing computational overhead. The hyperparameter selection details for the 10 clustering algorithms considered in our study are available in Section 2.3.
2.3 Hyperparameters Used with Clustering Algorithms
Table 1: Summary of hyperparameters and search spaces used for each clustering algorithm used in our experiments.
Algorithm Hyperparameter Range / Search Values
$k$-means n_init 100
$k$-medians n_init 100
Ward n_neighbors $[3,10]$, step 1
Agglomerative (AC) linkage $\{\text{complete, average, single}\}$
metric $\{\text{euclidean, l1, l2, cosine}\}$
n_neighbors $[3,10]$, step 1
DBSCAN eps $[0.1,1.5]$, step 0.1
HDBSCAN min_samples $[3,10]$, step 1
min_cluster_size $[5,25]$, step 5
OPTICS min_samples $[3,10]$, step 1
min_cluster_size $[0.1,0.5]$, step 0.05
xi $[0.05,0.25]$, step 0.05
Spectral Clustering (SC) eigen_solver $\{\text{arpack, lobpcg}\}$
affinity $\{\text{nearest\_neighbors, rbf}\}$
BIRCH threshold $[0.1,1.5]$, step 0.1
GMM covariance_type $\{\text{full, tied, diagonal, spherical}\}$
To ensure a fair and comprehensive comparison across clustering algorithms, we systematically explored a curated set of hyperparameter configurations for each of them, while preserving computational tractability. The selected configurations aim to balance representational diversity and practical feasibility in automated clustering scenarios. Table 1 summarizes the search spaces and parameters employed in our simulation study. For all configurations, parameters not explicitly mentioned in this table were maintained at their default values as defined in the scikit-learn framework. The applied strategy ensures a reasonable balance between optimization quality and computational complexity for each clustering algorithm considered.
@@ 2.4 Deep Model Architecture
It starts with a convolutional layer, followed by batch normalization and max pooling. Integrating a 2D convolutional layer, batch normalization, and max pooling is a standard and effective way of organizing an initial processing block in contemporary CNN architectures. After the convolutional layer, the architecture incorporates two consecutive residual blocks. Each of them can be characterized by the following equation:
@@
After the residual blocks, we implement a self-attention mechanism inspired by the transformer architecture from Vaswani et al. [31]. The attention block can be described by the following equation:
@@
where $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ are learned linear projections of the input features. The attention block uses the learned query, key, and value projections to compute a weighted sum of the input features, enabling the model to capture long-range dependencies within the feature maps and complement the local processing capabilities of the convolutional layers. The output from the attention block is subsequently flattened and processed through two fully connected layers, reaching the final layer that produces logits corresponding to the number of classes in the classification task.
Therefore, the proposed network architecture is designed to integrate local spatial pattern recognition (via CNNs), stable and hierarchical feature propagation (via residual connections), and global context modeling (via attention), offering a unified solution for learning discriminative representations from raw tabular data. As we will see in the next sections, an effective combination of these components ClustRecNet to deliver quality performance on complex clustering classification tasks.
2.5 Implementation Details
The described network architecture was chosen after extensive experimentation with multiple network variants, as it provided the best overall performance across our synthetic data repository. Moreover, the ablation study conducted on real-world data (see Section 3.2) showed that all components of the proposed network architecture are essential for generalizing across diverse clustering scenarios.